OpenStudy (anonymous):

alg2 help 7/3x + 2/x+ x/x-1

OpenStudy (alekos):

so are we simplifying?

OpenStudy (korosh23):

I simplified it a little bit.

OpenStudy (korosh23):

7/3x+6/3x+x/x-1

OpenStudy (korosh23):

For the last one, I am not sure what denominator we should use. Anyone ?

OpenStudy (anonymous):

how do I find the denominator

OpenStudy (korosh23):

That is what I am confused about ?

OpenStudy (misty1212):

\[\huge \frac{7}{3x}+\frac{2}{x}+\frac{x}{x-1}\]??

OpenStudy (misty1212):

common denominator would be \[3x(x-1)\] if that is the question

OpenStudy (alekos):

ladystar. send us a screen shot of the question

OpenStudy (anonymous):

misty has it but how do I find the denominator

OpenStudy (misty1212):

i wrote it it is\[\huge 3x(x-1)\]

OpenStudy (misty1212):

\[ \frac{7}{3x}+\frac{2}{x}+\frac{x}{x-1}\]\[\frac{7}{3x}\times \frac{x-1}{x-1}+\frac{2}{x}\times\frac{ 3(x-1)}{3(x-1)}+\frac{x}{x-1}\times \frac{3x}{3x}\]\]